OpenStudy (anonymous):

How do you solve: 1) 2^(x-4)+10=22 2) Log base 3 of 3 + log base 3 of x = 5 3) 3(5)^(2x-3)=6

OpenStudy (anonymous):

\[\bf 2^{x-4}=12\]Take log of both sides and solve for 'x'. \[\bf \log_3(3)+\log_3(x)=5 \implies \log_3(3 \times x)=5\]Now solve for 'x' from here. \[\bf 3(5)^{2x-3}=6 \implies 5^{2x-3}=3\]Take log of both sides and solve for 'x'. If you need any further help, tag me and I'll drop by. @Nekojellehfish

OpenStudy (anonymous):

@genius12 Ok, thanks!